OpenStudy (anonymous):

What is the value of the function y = 1.3x + 22 when x = −7? y =???

Nnesha (nnesha):

just substitute x value and then solve

OpenStudy (anonymous):

1.3+22=23.3 so Y=23.3

Nnesha (nnesha):

substitute -7 into x

OpenStudy (anonymous):

30.3

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

nope how did u plug x values show your work please so i can figure our ur mistake :-)

OpenStudy (anonymous):

1.3+22=23.3 then 23.3-(-7)

Nnesha (nnesha):

nope first you have to distribute parentheses

Nnesha (nnesha):

\[\huge\rm y = 1.3(-7) +22\] replaced x by -7 now distribute first and then add

OpenStudy (anonymous):

so 1.3*7+22?

Nnesha (nnesha):

look at the equation it says x = -7 not positive seven so substitute - 7

Nnesha (nnesha):

question** not equation :-)

OpenStudy (anonymous):

would I minus seven from the equation and then do that to the other side?.... if we are tring to get rid of it/

Nnesha (nnesha):

first multiply -7 times 1.3= ??

OpenStudy (anonymous):

-7*1.3=-9.1 and then -9.1+22?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

12.9

Nnesha (nnesha):

yep

Nnesha (nnesha):

:)

OpenStudy (anonymous):

thank you again

Nnesha (nnesha):

np :)